# Advancing Video Anomaly Detection: A Concise Review and a New Dataset

### v2

@@ Appendix A Dataset collection process
We have obtained permission from YouTube, Itemfix, and Bilibili to use their video data for non-commercial purposes. To facilitate the data collection, we use a script to automatically download relevant videos from these platforms and trim them into short clips. We also release these scripts along with our dataset for interested researchers.
Real-world scenarios present greater challenges due to diverse weather and lighting conditions, as well as unpredictable objects that can impact detection, such as dynamic electronic advertising screens or moving car headlights during nighttime in video surveillance. Our dataset considers these variations. Our test set covers a broad spectrum of abnormal events observed in diverse scenarios, encompassing both human-related and non-human-related anomalies, such as fire incidents, fighting, shooting, traffic accidents, falls, and other anomalies. Our test set exclusively contains video frame-level annotations. We apply a rigorous filtering process to ensure the quality and appropriateness of the collected videos. The following criteria are used to remove unsuitable videos:
@@ Appendix C Dataset usage and maintenance
The use of this dataset is governed by the following terms and conditions:
@@
Our project website can be found here.
@@ Appendix D Scenario-adaptive anomaly detection framework
Few-shot multi-scenario multi-view learning. Using a few-shot learning framework for anomaly detection is not an entirely new concept. One of the most widely recognized methods in this domain is the Few-shot Scene-adaptive Anomaly Detection (FSAD) model [36]. This model leverages the meta-learning framework [19] to train a model using video data collected from various camera views within the same scenario333In this paper, the term ‘scenario’ is employed to signify different contexts such as retail, manufacturing, the education sector, smart cities, and others. In the context of different camera views, existing literature sometimes uses the term ‘scene’ to refer to the camera scene from a specific camera viewpoint., such as a university street. The model is then fine-tuned on a different camera viewpoint within the university site. Although the trained model can be adapted to novel viewpoints, still its adaptability is confined to the university scenario. To overcome this limitation, we introduce the Scenario-Adaptive Anomaly Detection (SA2D) method. Diverging from existing few-shot anomaly detection models [36], our approach: (i) builds on and extends one-scenario multi-view to multi-scenario multi-view learning problem, and (ii) broadens the scope of test cases from multiple camera views to encompass novel scenarios, as well as multiple camera views per scenario. Our model stands out as a more advanced and versatile version in comparison to existing anomaly detection models. Fig. 6 shows a comparison between the existing few-shot anomaly detection model and our SA2D model.
@@ Appendix F Further evaluations
Performance comparison by anomaly type. Table 7 presents the frame-level Micro AUC and Average Precision (AP) performance, in percentage, for three weakly-supervised methods (RTFM [76], MGFN [13], and UR-DMU [102]) pretrained on either UCF-Crime or our MSAD training set. These methods are applied to our MSAD test set across various anomaly types (a total of 11 main anomaly types, as shown in Fig. 3(a)). We use I3D as the backbone for all methods in these experiments. We observe that methods pretrained on our MSAD dataset generally achieve better performance on anomalies such as fighting, fire, object falling, shooting, traffic accidents, and water incidents. Overall, using our dataset leads to better performance (see the overall performance column) compared to using UCF-Crime. This indicates that existing datasets do not focus sufficiently on non-human-related anomalies, and our dataset helps bridge this gap. However, for some anomaly types, such as explosions, people falling, and vandalism, using the UCF-Crime dataset achieves slightly better performance. The potential reasons for this are: (i) the I3D backbone encoder’s difficulty in capturing sudden motions, as it was originally designed for action recognition, and (ii) limitations of our dataset, which suggests the need for further collection and augmentation of video samples to improve model performance. We also note that there is no single best performer for all anomaly types, indicating that our dataset can support the VAD community in exploring and developing more robust anomaly detection algorithms.
@@ Appendix G A review of advances in datasets
From single-view to multi-view. Existing datasets fall into two main categories: single-view, such as UCSD Ped [79] and CUHK Avenue [35], and multi-view like ShanghaiTech [37] and UBnormal [2]. UCSD Ped [79] and CUHK Avenue [35] primarily originate from campus surveillance videos. ShanghaiTech [37], a pioneering multi-view dataset, has served as a benchmark for various methods. However, akin to other university datasets, it lacks object diversity and deviates from real-world scenarios regarding environment variations. UCF-Crime [72] was the first real-world dataset with multiple views. Despite encompassing diverse anomaly types, including abuse, explosions, stealing, and fighting, its quality is suboptimal due to monochrome video footage, low resolution, moving cameras, and redundancy. Still, most existing datasets focus on human-related anomalies.
From single-scenario to multi-scenario. Single-scenario datasets [6, 35, 79, 37, 63] are typically sourced from specific locations with one or multiple camera viewpoints, such as universities or urban streets, capturing routine activities and occasional anomalies. Early datasets, such as UCSD Ped [79] and CUHK Avenue [35], were collected from a single camera viewpoint at a university. Although single-scenario datasets are valuable, they lack diversity in scenarios, each with distinct backgrounds and anomalies. For instance, a parking lot typically has less movement compared to a crowded street, and a shopping mall is busier than a university avenue. Training a model exclusively on a single scenario limits its ability to generalize, as it only learns the specific features of that scenario and performs poorly when applied to new scenes [36]. To address the limited real-world diversity, multi-scenario datasets, such as UCF-Crime [72] and CUVA [14], have been collected from online video platforms (e.g., YouTube), offering a wide range of scenarios. These variations in scenarios, with diverse weather and lighting conditions, greatly enhance the robustness and adaptability of anomaly detection models for real-world applications. UCF-Crime [72] has become the largest and most popular benchmark for weakly-supervised methods. The CUVA [14] dataset is another large-scale benchmark focused on the causation of video anomalies and is the first to include high-quality text descriptions of video content. This is designed to aid video large language models (VLMs) in understanding the cause and effect of anomalies. However, unlike UCF-Crime [72], which exclusively collects surveillance videos from websites, CUVA sources its videos from news outlets and handheld cameras. This results in significant viewpoint shifts that may affect performance in surveillance.
@@ Appendix H Limitation and future work
Although our dataset covers a wide range of scenarios and anomaly types, there are still more scenarios and anomaly types we need to consider for real-world applications. Some anomaly types, such as uncontrolled animals and sudden crowd gatherings, are still missing from our dataset. Our future work will focus on collecting more scenarios and anomaly types.
Videos for industrial scenarios, like adherence to safety protocols such as wearing hard hats or operating equipment safely, are difficult to obtain due to privacy issues. We will explore alternative solutions for these anomalies. Moreover, we will continue refining our dataset for existing anomalies, such as no right turns at specific crossings, for traffic-related anomalies.
Additionally, we will provide more modalities to further expand our dataset, such as audio, video captions, and human skeletons. Given recent advances, there is a greater focus on reusable large-scale pretrained models and multi-modal fusion for robust VAD.

### v3

@@ Appendix A Dataset collection process
We have obtained permission from YouTube, Itemfix, and Bilibili to use their video data for non-commercial purposes. To facilitate the data collection, we use a script to automatically download relevant videos from these platforms and trim them into short clips. We also release these scripts along with our dataset for interested researchers. In our collection process, most videos are collected from YouTube (717 videos), and only a few videos are collected from BiliBili and ItemFix (3 videos). We ensure that all the videos are selected and used in accordance with the Fair Dealing provisions of the Australian Copyright Act 1968. Our approach is designed to adhere strictly to these legal frameworks, ensuring that the use of copyrighted material is only limited to research purposes. This careful consideration helps to protect the rights of the original content creators while allowing us to utilize the material in a legally compliant manner.
Real-world scenarios present greater challenges due to diverse weather and lighting conditions, as well as unpredictable objects that can impact detection, such as dynamic electronic advertising screens or moving car headlights during nighttime in video surveillance. Our dataset considers these variations. Our test set covers a broad spectrum of abnormal events observed in diverse scenarios, including both human-related and non-human-related anomalies, such as fire incidents, fighting, shooting, traffic accidents, falls, and other anomalies. Our test set exclusively contains video frame-level annotations. We apply a rigorous filtering process to ensure the quality and appropriateness of the collected videos. The following criteria are used to remove unsuitable videos:
@@
While we acknowledge that surveillance videos can raise privacy and violence concerns, the potential risks are magnified when anomalies occur. A real-world anomaly detection dataset serves as a necessary complement, despite the inherent drawbacks of surveillance. However, it is crucial to implement appropriate regulations to minimize the potential negative impact of datasets. Our MSAD dataset is intended exclusively for academic research. Researchers who wish to access the dataset must complete the online form and agree to the terms and conditions.
Identifiable information. To eliminate any identifiable information, we consider the following options. (i) Face blurring: Since in surveillance video footage, the facial regions of individuals are often small and blurry, or the camera angles typically capture side or rear views, it is challenging to extract identifiable information about individuals. To eliminate any identifiable information that might be visible, we apply an automatic blurring script444https://github.com/ORB-HD/deface to anonymize all the faces and car licenses in the videos of the MSAD dataset, then we check the whole dataset manually to ensure that all the identity information is almost eliminated. The blurred videos have limited impact to the detection performance while reducing the invasion of privacy. We provide the blurred version of our MSAD dataset. (ii) Extracted features: We have also released the extracted video features with different backbones, e.g., I3D, SwinTransformer, etc., to the research community to promote privacy preservation. We have noticed that some weakly-supervised methods in recent years use extracted video features as inputs [83, 14, 13, 108]. Training and evaluation with extracted video features is lightweight while protecting privacy. (iii) Deepfake technique: We also explore using deepfake techniques to remove identifiable features from the original videos, aiming to preserve characteristics such as age and gender while minimizing information loss compared to face blurring. However, we encounter challenges due to the diverse resolutions and camera angles, which make deepfake difficult. We plan to further investigate this approach in future work.
Bias. To ensure diversity in our dataset, we have extensively collected videos from various regions and countries. The purpose of collecting videos from diverse regions and countries is to advance anomaly detection technologies globally, benefiting communities worldwide. We conduct searches on platforms such as YouTube using text queries that combine various anomaly types (e.g., fighting, fire, people falling, etc.) with diverse region and country-specific terms. To maximize the diversity of videos retrieved from around the world, we also utilize queries in multiple languages using translators. The collected videos represent a broad spectrum of ages, genders, ethnicities, and scenarios, which helps to reduce potential biases. By incorporating such a diverse array of content, our goal is to create a balanced and representative dataset that accurately mirrors the complexity and diversity of real-world situations.
@@ Appendix C Dataset usage and maintenance
The use of our dataset is governed by the following terms and conditions:
@@
To address concerns about deprecated videos, we will regularly monitor the sources of our dataset to ensure that the videos or data links remain valid. This process may involve automated scripts that periodically verify online links and alert us if any sources become inaccessible. If the original video source becomes unavailable, we will seek alternative sources or mirrors that host the same content. Maintaining a list of alternative sources will help preserve the dataset’s integrity. However, if no alternative sources are found, we may consider removing the unavailable content from our dataset to ensure that it remains up-to-date and reliable.
Additionally, we will update the experimental results on our dataset website if any videos become unavailable. We will also document the missing videos for transparency and fairness in future research comparisons. We will provide a detailed datasheet to facilitate a more comprehensive assessment of the benchmark and baselines. We have provided our script for downloading the videos, as well as extracted features from backbone networks such as I3D and SwinTransformer. We will also release our full evaluation framework to allow interested researchers to further explore our multi-scenario datasets in the project website. We believe that our efforts provide a solid foundation for future research on real-world anomaly detection, ultimately benefiting our community.
@@ Appendix D Scenario-adaptive anomaly detection framework
Few-shot multi-scenario multi-view learning. Using a few-shot learning framework for anomaly detection is not an entirely new concept. One of the most widely recognized methods in this domain is the Few-shot Scene-adaptive Anomaly Detection (FSAD) model [42]. This model leverages the meta-learning framework [22] to train a model using video data collected from various camera views within the same scenario555In this paper, the term ‘scenario’ is employed to signify different contexts such as retail, manufacturing, the education sector, smart cities, and others. In the context of different camera views, existing literature sometimes uses the term ‘scene’ to refer to the camera scene from a specific camera viewpoint., such as a university street. The model is then fine-tuned on a different camera viewpoint within the university site. Although the trained model can be adapted to novel viewpoints, still its adaptability is confined to the university scenario. To overcome this limitation, we introduce the Scenario-Adaptive Anomaly Detection (SA2D) method. Diverging from existing few-shot anomaly detection models [42], our approach: (i) builds on and extends one-scenario multi-view to multi-scenario multi-view learning problem, and (ii) broadens the scope of test cases from multiple camera views to encompass novel scenarios, as well as multiple camera views per scenario. Our model stands out as a more advanced and versatile version in comparison to existing anomaly detection models. Fig. 6 shows a comparison between the existing few-shot anomaly detection model and our SA2D model.
@@ Appendix F Further evaluations
Performance comparison by anomaly type. Table 7 presents the frame-level Micro AUC and Average Precision (AP) performance, in percentage, for three weakly-supervised methods (RTFM [83], MGFN [14], and UR-DMU [117]) pretrained on either UCF-Crime or our MSAD training set. These methods are applied to our MSAD test set across various anomaly types (a total of 11 main anomaly types, as shown in Fig. 3(a)). We use I3D as the backbone for all methods in these experiments. We observe that methods pretrained on our MSAD dataset generally achieve better performance on anomalies such as fighting, fire, object falling, shooting, traffic accidents, and water incidents. Overall, using our dataset leads to better performance (see the overall performance column) compared to using UCF-Crime. This indicates that existing datasets do not focus sufficiently on non-human-related anomalies, and our dataset helps bridge this gap. However, for some anomaly types, such as explosions, people falling, and vandalism, using the UCF-Crime dataset achieves slightly better performance. The potential reasons for this are: (i) the I3D backbone encoder’s difficulty in capturing sudden motions, as it is originally designed for action recognition, and (ii) limitations of our dataset, which suggests the need for further collection and augmentation of video samples to improve model performance. We also note that there is no single best performer for all anomaly types, indicating that our dataset can support the VAD community in exploring and developing more robust anomaly detection algorithms.
@@
Content differences between normal and abnormal videos. In fact, most existing anomalous videos exhibit significant visual differences from normal videos in the test set, in some existing datasets such as ShanghaiTech. To mitigate this, when collecting our dataset, we split long videos from the same camera viewpoint into shorter clips to ensure that these videos are not influenced by visual differences between normal and abnormal scenes. This approach allows our dataset to be used for exploring real anomalies rather than global frame textures / contexts. As shown in Table 7 and 8, our dataset can evaluate model performance on different anomaly types independent of scenarios/contexts and on different scenarios independent of anomaly types. For example, the Object Falling anomalies happen under diverse scenes including the front door, office, road, etc. However, the model also gained good performance (above 90% AUC score), it can be concluded that the model focuses more on actions rather than spurious features.
Below we provide more insights on our new dataset. (i) Diverse video scenarios: Our training and testing videos are sourced from a wide range of scenarios, all featuring complex environments such as busy supermarkets/streets, empty parking lots/stores, and footage captured during different times of the day (morning/night). This diversity makes it challenging for the model to learn from trivial spurious features like global frame textures. (ii) Model performance in different scenarios: Tables 7 and 8 show that the current models struggle in certain scenarios (low average precision) and do not perform well across all types of anomalies. For instance, the detection of anomalies like object falling achieves good results (over 90% AUC) in diverse scenes, including the front door, office, and road. This indicates that the model relies more on motion rather than scene characteristics for judgment. However, in some scenarios, like a pedestrian street, while the mean AUC across three methods reaches 89.1%, the AP is only 27.5%. This highlights the significant challenge posed by the complex scenes (e.g., crowded streets) in our dataset to current detection methods. (iii) Feature magnitude analysis: We analyze the feature magnitude extracted from our dataset using different backbones (e.g., I3D, SwinTransformer) for each snippet (16 frames). In previous datasets (e.g., UCSD Ped, CUHK Avenue, UCF-Crime), the snippet-level feature magnitude increases significantly during anomalies, while it remains low during normal situations due to simpler scenes. However, in our dataset, this assumption, using feature magnitude to differentiate normal from abnormal events, does not hold, as normal situations in complex environments (e.g., busy streets) also exhibit significant and noisy changes in feature magnitude. We aim to use our MSAD dataset to advance anomaly detection in more challenging scenarios in future work.
Concerns on spurious correlations in anomalous event detection. We acknowledge the potential for a model to learn spurious correlations related to gender, race, or location when training on our dataset, especially in action classes like assault, fighting, and robbery, where faces and locations may be visible. We recognize the risk that a black-box video anomaly detection model could inadvertently learn such correlations, leading to biased or unfair outcomes. To address this, we propose the following recommendations and safeguards: (i) implementing additional anonymization techniques in the dataset, such as blurring or masking faces and identifiable features in the training data, with our blurred version of MSAD provided for researchers; (ii) using balanced sampling strategies and data augmentation techniques to ensure diverse representation across demographics and locations, minimizing the risk of unintended associations; (iii) incorporating post-training bias detection methods to identify learned spurious correlations, applying techniques such as fairness-aware learning [36] to mitigate these biases during training and evaluation; (iv) encouraging the use of explainable AI methods [16, 17] in anomaly detection to provide insights into the model’s decision features, promoting transparency to identify and address spurious correlations early; and (v) proposing continuous monitoring and evaluation of the model on diverse, representative test sets to ensure it generalizes well and avoids biased behavior in different contexts. In future, we plan to incorporate these recommendations more explicitly and explore further safeguards to ensure the ethical use of our dataset in training video anomaly detection models.
@@ Appendix G A review of advances in datasets
From single-view to multi-view. Existing datasets fall into two main categories: single-view, such as UCSD Ped [86] and CUHK Avenue [41], and multi-view like ShanghaiTech [43] and UBnormal [2]. UCSD Ped [86] and CUHK Avenue [41] primarily originate from campus surveillance videos. ShanghaiTech [43], a pioneering multi-view dataset, has served as a benchmark for various methods. However, akin to other university datasets, it lacks object diversity and deviates from real-world scenarios regarding environment variations. UCF-Crime [79] is the first real-world dataset with multiple views. Despite encompassing diverse anomaly types, including abuse, explosions, stealing, and fighting, its quality is suboptimal due to monochrome video footage, low resolution, moving cameras, and redundancy. Still, most existing datasets focus on human-related anomalies.
From single-scenario to multi-scenario. Single-scenario datasets [6, 41, 86, 43, 70] are typically sourced from specific locations with one or multiple camera viewpoints, such as universities or urban streets, capturing routine activities and occasional anomalies. Early datasets, such as UCSD Ped [86] and CUHK Avenue [41], are collected from a single camera viewpoint at a university. Although single-scenario datasets are valuable, they lack diversity in scenarios, each with distinct backgrounds and anomalies. For instance, a parking lot typically has less movement compared to a crowded street, and a shopping mall is busier than a university avenue. Training a model exclusively on a single scenario limits its ability to generalize, as it only learns the specific features of that scenario and performs poorly when applied to new scenes [42]. To address the limited real-world diversity, multi-scenario datasets, such as UCF-Crime [79] and CUVA [17], have been collected from online video platforms (e.g., YouTube), offering a wide range of scenarios. These variations in scenarios, with diverse weather and lighting conditions, greatly enhance the robustness and adaptability of anomaly detection models for real-world applications. UCF-Crime [79] has become the largest and most popular benchmark for weakly-supervised methods. The CUVA [17] dataset is another large-scale benchmark focused on the causation of video anomalies and is the first to include high-quality text descriptions of video content. This is designed to aid video large language models (VLMs) in understanding the cause and effect of anomalies. However, unlike UCF-Crime [79], which exclusively collects surveillance videos from websites, CUVA sources its videos from news outlets and handheld cameras. This results in significant viewpoint shifts that may affect performance in surveillance.
@@ Appendix H Limitation and future work
For evaluation criteria, we use frame-level AUC and average precision for temporal dimension, however, we also note that some benchmarks (like UCSD Ped, CUHK, etc.) have pixel-level annotations for spatial anomaly localizations. This kind of annotation is time-consuming and label intensive for large benchmarks. Looking ahead, we plan to enhance our dataset to support more comprehensive evaluations. This may include incorporating additional textual descriptions or adding bounding box annotations, which would enable our dataset to be used for spatial anomaly localization and causal reasoning. By extending our dataset in this manner, we aim to facilitate more advanced research in anomaly detection. Furthermore, we intend to integrate spatio-temporal localization-based metrics, such as region- and time-based detection scores, to further advance video anomaly detection by exploring fine-grained anomaly regions within videos. We plan to include this evaluation metric in a future version of our dataset.
We focus on weakly-supervised learning methods in this work. We will consider both parametric and non-parametric models in the future scope of our work. While recent methods have largely focused on parametric approaches, it is inspiring that earlier research works use non-parametric methods for anomaly detection. These non-parametric techniques often involve analyzing optical flow to detect motion changes [27, 51], offering an interpretable perspective on anomaly perception. In our future work, we plan to draw inspiration from these methods to evaluate the disorder in motion changes.
Although our dataset covers a wide range of scenarios and anomaly types, there are still more scenarios and anomaly types we need to consider for real-world applications. Some anomaly types, such as uncontrolled animals and sudden crowd gatherings, are still missing from our dataset. Our future work will focus on collecting more scenarios and anomaly types. Videos for industrial scenarios, like adherence to safety protocols such as wearing hard hats or operating equipment safely, are difficult to obtain due to privacy issues. We will explore alternative solutions for these anomalies. Moreover, we will continue refining our dataset for existing anomalies, such as no right turns at specific crossings, for traffic-related anomalies. Additionally, we will provide more modalities to further expand our dataset, such as audio, video captions, and human skeletons. Given recent advances, there is a greater focus on reusable large-scale pretrained models and multi-modal fusion for robust VAD.